# Wonder3D++: Cross-domain Diffusion for High-fidelity 3D Generation from a Single Image

### v1

@@ IV-A3 Cross-domain Attention
Given a latent embedding $Z$, the original self-attention mechanism [69] can be defined as:
@@ IV-A4 Multi-view Cross-domain Transformer Block
The prior 2D diffusion models [54, 36] generate each image separately, so that the resulting images are not geometrically and visually consistent across different views. To enhance consistency among different views, similar to prior works such as SyncDreamer [38] and MVDream [60], we utilize attention mechanism to facilitate information propagation across different views, implicitly encoding multi-view dependencies (as illustrated in Figure 3). This is achieved by extending the original self-attention layers to be global-aware, allowing connections to other views within the attention layers. Keys and values from different views are connected to each other to facilitate the exchange of information. By sharing information across different views and domains within the attention layers, the model perceives cross-domain multi-view correlation and is capable of generating consistent multi-view color images and normal maps.
@@ IV-B1 Geometric Initialization
Explicit mesh optimization methods typically require a mesh with approximate topology for initialization. Following [79], we first use normal map integration from the front and back views to obtain two corresponding depth maps. These depth maps are then projected into 3D space to create two point cloud parts with normal, which are combined into an initialized mesh using the Poisson reconstruction [27] method. Unlike the commonly used sphere initialization, this approach not only speeds up optimization but also enables our method to capture relatively complex shapes by providing an approximate topological structure.
@@
Figure 4: The illustration of our geometric initialization strategy. We employ the Poisson reconstruction [27] method for geometric initialization using normal maps from front and back views, while estimated depth maps are used to detect and correct potential initialization errors.
@@ IV-B2 Coarse Reconstruction
Realfusion [46] 0.0819 $\pm$ 0.0010 0.2741 $\pm$ 0.0092
Magic123 [53] 0.0516 $\pm$ 0.0021 0.4528 $\pm$ 0.0183
One-2-3-45 [35] 0.0629 $\pm$ 0.0009 0.4086 $\pm$ 0.0125
Point-E [49] 0.0426 $\pm$ 0.0011 0.2875 $\pm$ 0.0071
Shap-E [25] 0.0436 $\pm$ 0.0018 0.3584 $\pm$ 0.0088
Zero123 [36] 0.0339 $\pm$ 0.0015 0.5035 $\pm$ 0.0084
SyncDreamer [38] 0.0261 $\pm$ 0.0062 0.5421 $\pm$ 0.0079
OpenLRM [21] 0.0255 $\pm$ 0.0027 0.5452 $\pm$ 0.0064
InstantMesh [82] 0.0224 $\pm$ 0.0015 0.5353 $\pm$ 0.0084
CRM [75] 0.0220 $\pm$ 0.0009 0.5412 $\pm$ 0.0059
Unique3D [75] 0.0238 $\pm$ 0.0023 0.5134 $\pm$ 0.0112
Era3D  [32] 0.0239 $\pm$ 0.0026 0.5340 $\pm$ 0.0075
Wonder3D  [41] 0.0222 $\pm$ 0.0011 0.5521 $\pm$ 0.0066
@@ IV-B3 Iterative Refinement
The module uses the ControlNet [91] architecture to fully leverage prior information from stable diffusion. We treat the ControlNet as a trainable module while keeping the pre-trained stable diffusion weights fixed. To maintain consistency between the enhanced normal maps and color images, we integrate cross-domain attention into the ControlNet. Low-quality RGB images and normal maps are input into the enhancement module, producing corresponding high-quality outputs with 2x resolution upsampling and enhancement. Additionally, the original input image is processed through an IP-Adapter [84] to incorporate high-level textual information.
TABLE II: Quantitative comparison for novel view synthesis. We report PSNR, SSIM [76], LPIPS [92] on the GSO [14] dataset, where $\pm$ indicates the standard deviation.
@@
Realfusion [46] 15.26 $\pm$ 0.24 0.722 $\pm$ 0.017 0.283 $\pm$ 0.005
Zero123 [36] 18.93 $\pm$ 0.22 0.779 $\pm$ 0.015 0.166 $\pm$ 0.005
SyncDreamer [38] 20.05 $\pm$ 0.22 0.798 $\pm$ 0.013 0.146 $\pm$ 0.004
SV3D [70] 21.26 $\pm$ 0.26 0.880 $\pm$ 0.015 0.080 $\pm$ 0.005
Unique3D [79] 21.71 $\pm$ 0.19 0.913 $\pm$ 0.012 0.083 $\pm$ 0.003
Era3D [32] 22.49 $\pm$ 0.24 0.916 $\pm$ 0.011 0.069 $\pm$ 0.004
Wonder3D [41] 22.92 $\pm$ 0.22 0.919 $\pm$ 0.009 0.063 $\pm$ 0.002
@@ V-A Implementation Details
We train our model on a high-quality single object subset of the Objaverse dataset [10], which comprises approximately 100k objects following a cleanup process. To support the model in learning consistent representations across diverse camera settings, we create the rendered augmented multi-view dataset. We first normalized each object to be centered and on a unit scale. Then we render normal maps and color images from six views, including the front, back, left, right, front-right, and front-left views, using Blenderproc [12] under both orthogonal and perspective camera settings.
@@ V-B Baselines
We adopt a comprehensive set of baseline methods across three key categories: SDS-based methods, multi-view-based methods, and feed-forward methods. For SDS-based methods, RealFusion [46] and Magic123 [53] leverage Score Distillation Sampling (SDS) loss for guiding single-view 3D reconstruction from an input image. In the MV-based category, Zero123 [36] generates novel views from arbitrary perspectives, which can be further combined with SDS loss [52] to perform 3D reconstruction. Additionally, SyncDreamer [38], Era3D [32], and Wonder3D [41] produce multi-view consistent images, using the Neus [72] representation to reconstruct 3D geometry. In contrast, Unique3D [79] directly optimizes mesh vertices for a reconstruction approach. For feed-forward methods, Point-E [49] and Shap-E [25] are 3D generative models trained on a large internal OpenAI 3D dataset, both of which are able to convert a single-view image into a point cloud or an implicit representation. One-2-3-45 [35], CRM [75] and InstantMesh [82] first generate multi-view results and then leverage a sparse-view feed-forward reconstruction model to deliver a textured mesh using the representation of triplane or implicit SDF representation.
@@ V-C Evaluation Protocol
Evaluation Datasets. Following prior research [36, 38], we adopt the Google Scanned Object dataset [14] for our evaluation, which includes a wide variety of common everyday objects. Our evaluation dataset aligns with that of SyncDreamer [38], comprising 30 objects spanning from everyday items to animals. For each object in the evaluation set, we render an image with a size of 256×256, which serves as the input. Additionally, we include in-the-wild images with diverse styles, collected from the internet or generated using text-to-image models, in our evaluation.
Metrics. To evaluate the quality of the single-view reconstructions, we adopt two metrics: Chamfer Distances (CD) and Volume IoU between ground-truth shapes and reconstructed shapes on both orthogonal and perspective settings of our method. Since different methods adopt various canonical systems, we first align the generated shapes to the ground-truth shapes before calculating the two metrics. Moreover, we adopt the metrics PSNR, SSIM [76] and LPIPS [92] for evaluating the generated color images.
@@ V-D Single View Reconstruction
We compare the quality of the reconstructions with SOTA methods, the quantitative results are summarized in Table I, and the qualitative comparisons are presented in Fig. 6. Rather than modeling multi-view images and normal maps distribution jointly, Unique3D [79] sequentially generate multi-view colors and normal maps. This exacerbates the geometric inconsistency since the stage 2 model has a huge domain gap in the inputs during training and in the inference period(see Section V-F for a more detailed discussion). As a result, vertices are updated by unstable and incorrect gradients during the optimization process, destabilizing the process and ultimately leading to a reconstructed object with incorrect shapes. Leveraging the generalization performance of the multi-view diffusion model and sparse-view reconstruction model, InstantMesh [82] produces results with reasonable overall geometry. However, due to the limited resolution of the triplane representation, the generated geometry and textures lack sufficient detail and exhibit artifacts aligned perpendicular to the coordinate axes. Wonder3D [41] takes relatively low-resolution multi-view images and normal maps to reconstruct 3D objects using implicit SDF representation without an enhancement process, it also struggles to obtain fine geometric and texture details.
Thanks to our multi-stage training scheme and cascaded 3D mesh extraction, our method produces textured meshes with a well-defined overall shape and superior geometric and textural quality compared to baseline approaches. Furthermore, geometric quantization comparisons on the GSO dataset [14] highlight the generalization ability and effectiveness of our method (see Table  I).
@@ V-E Novel View Synthesis
We evaluate the quality of novel view synthesis for different multi-view-based methods. The quantitative results are presented in Table II, and the qualitative results can be found in Figure 5. SyncDreamer [36] introduces a volume attention scheme to enhance the consistency of multi-view images, their model is sensitive to the elevation degrees of the input images and tends to produce unreasonable results. Unique3D [79] stitching together the multiple views into a single image to generate multi-view results. This method severely limits the number of viewpoints it can produce (up to four). Moreover, generating images and normal maps in a sequential manner limits the generalizability and cross-view consistency of the model. Era3D [32] proposes row-wise attention to ensure multi-view consistency and estimates the input camera focal length to normalize the output results into an orthogonal coordinate system. Although it generates results with a higher resolution, it is difficult to stably output reasonable multi-view results for in-the-wild images.
@@ V-F Discussions
Figure 11: The qualitative comparisons of reconstruction method between the geometry fusion algorithm proposed by Wonder3D [41] and cascaded 3D mesh extraction method. Our approach has significant advantages in both geometric and textural details.
@@ Acknowledgment
This work was supported by Project NSFC 62501261.
@@ Appendix A Coordinate System
In practice, the target object is assumed to be placed along the gravity direction. 1) Canonical coordinate system. Some prior works (e.g. MVDream [60] and SyncDreamer [38]) adopt a shared canonical system for all objects, whose axis $Z_{c}$ shares the same direction with gravity (Fig. 15 (a)). 2) Input view related system. Wonder3D and Wonder3D++ adopt an independent coordinate system for each object that is related to the input view. Its $Z_{v}$ and $X_{v}$ axes are aligned with the UV dimension of 2D input image space, and its $Y_{v}$ axis is vertical to the 2D image plane and passes through the center of ROI (Region of Interests) (Fig. 15 (b)). 3) Camera poses. Wonder3D and Wonder3D++ outputs 6 views $\{v_{i},i=0,...,5\}$ that are sampled at the $X_{v}OY_{v}$ plane of the input-view related system with a fixed radius, where the front view $v_{0}$ is initialized as input view and the other views are sampled with pre-defined azimuth degrees (see Fig. 15 (b)).

### v2

@@ IV-A3 Cross-domain Attention
Given a latent embedding $Z$, the original self-attention mechanism [vaswani2017attention] can be defined as:
@@ IV-A4 Multi-view Cross-domain Transformer Block
The prior 2D diffusion models [radford2021learning, liu2023zero] generate each image separately, so that the resulting images are not geometrically and visually consistent across different views. To enhance consistency among different views, similar to prior works such as SyncDreamer [liu2023syncdreamer] and MVDream [shi2023mvdream], we utilize attention mechanism to facilitate information propagation across different views, implicitly encoding multi-view dependencies (as illustrated in Figure 3). This is achieved by extending the original self-attention layers to be global-aware, allowing connections to other views within the attention layers. Keys and values from different views are connected to each other to facilitate the exchange of information. By sharing information across different views and domains within the attention layers, the model perceives cross-domain multi-view correlation and is capable of generating consistent multi-view color images and normal maps.
@@ IV-B1 Geometric Initialization
Explicit mesh optimization methods typically require a mesh with approximate topology for initialization. Following [wu2024unique3d], we first use normal map integration from the front and back views to obtain two corresponding depth maps. These depth maps are then projected into 3D space to create two point cloud parts with normal, which are combined into an initialized mesh using the Poisson reconstruction [kazhdan2006poisson] method. Unlike the commonly used sphere initialization, this approach not only speeds up optimization but also enables our method to capture relatively complex shapes by providing an approximate topological structure.
@@
Figure 4: The illustration of our geometric initialization strategy. We employ the Poisson reconstruction [kazhdan2006poisson] method for geometric initialization using normal maps from front and back views, while estimated depth maps are used to detect and correct potential initialization errors.
@@ IV-B2 Coarse Reconstruction
Realfusion [melas2023realfusion] 0.0819 $\pm$ 0.0010 0.2741 $\pm$ 0.0092
Magic123 [qian2023magic123] 0.0516 $\pm$ 0.0021 0.4528 $\pm$ 0.0183
One-2-3-45 [liu2023one] 0.0629 $\pm$ 0.0009 0.4086 $\pm$ 0.0125
Point-E [nichol2022point] 0.0426 $\pm$ 0.0011 0.2875 $\pm$ 0.0071
Shap-E [jun2023shap] 0.0436 $\pm$ 0.0018 0.3584 $\pm$ 0.0088
Zero123 [liu2023zero] 0.0339 $\pm$ 0.0015 0.5035 $\pm$ 0.0084
SyncDreamer [liu2023syncdreamer] 0.0261 $\pm$ 0.0062 0.5421 $\pm$ 0.0079
OpenLRM [hong2023lrm] 0.0255 $\pm$ 0.0027 0.5452 $\pm$ 0.0064
InstantMesh [xu2024instantmesh] 0.0224 $\pm$ 0.0015 0.5353 $\pm$ 0.0084
CRM [wang2024crm] 0.0220 $\pm$ 0.0009 0.5412 $\pm$ 0.0059
Unique3D [wang2024crm] 0.0238 $\pm$ 0.0023 0.5134 $\pm$ 0.0112
Era3D  [li2024era3d] 0.0239 $\pm$ 0.0026 0.5340 $\pm$ 0.0075
Wonder3D  [long2024wonder3d] 0.0222 $\pm$ 0.0011 0.5521 $\pm$ 0.0066
@@ IV-B3 Iterative Refinement
The module uses the ControlNet [zhang2023adding] architecture to fully leverage prior information from stable diffusion. We treat the ControlNet as a trainable module while keeping the pre-trained stable diffusion weights fixed. To maintain consistency between the enhanced normal maps and color images, we integrate cross-domain attention into the ControlNet. Low-quality RGB images and normal maps are input into the enhancement module, producing corresponding high-quality outputs with 2x resolution upsampling and enhancement. Additionally, the original input image is processed through an IP-Adapter [ye2023ip] to incorporate high-level textual information.
TABLE II: Quantitative comparison for novel view synthesis. We report PSNR, SSIM [wang2004image], LPIPS [zhang2018unreasonable] on the GSO [downs2022google] dataset, where $\pm$ indicates the standard deviation.
@@
Realfusion [melas2023realfusion] 15.26 $\pm$ 0.24 0.722 $\pm$ 0.017 0.283 $\pm$ 0.005
Zero123 [liu2023zero] 18.93 $\pm$ 0.22 0.779 $\pm$ 0.015 0.166 $\pm$ 0.005
SyncDreamer [liu2023syncdreamer] 20.05 $\pm$ 0.22 0.798 $\pm$ 0.013 0.146 $\pm$ 0.004
SV3D [voleti2024sv3d] 21.26 $\pm$ 0.26 0.880 $\pm$ 0.015 0.080 $\pm$ 0.005
Unique3D [wu2024unique3d] 21.71 $\pm$ 0.19 0.913 $\pm$ 0.012 0.083 $\pm$ 0.003
Era3D [li2024era3d] 22.49 $\pm$ 0.24 0.916 $\pm$ 0.011 0.069 $\pm$ 0.004
Wonder3D [long2024wonder3d] 22.92 $\pm$ 0.22 0.919 $\pm$ 0.009 0.063 $\pm$ 0.002
@@ V-A Implementation Details
We train our model on a high-quality single object subset of the Objaverse dataset [deitke2023objaverse], which comprises approximately 100k objects following a cleanup process. To support the model in learning consistent representations across diverse camera settings, we create the rendered augmented multi-view dataset. We first normalized each object to be centered and on a unit scale. Then we render normal maps and color images from six views, including the front, back, left, right, front-right, and front-left views, using Blenderproc [Denninger2023] under both orthogonal and perspective camera settings.
@@ V-B Baselines
We adopt a comprehensive set of baseline methods across three key categories: SDS-based methods, multi-view-based methods, and feed-forward methods. For SDS-based methods, RealFusion [melas2023realfusion] and Magic123 [qian2023magic123] leverage Score Distillation Sampling (SDS) loss for guiding single-view 3D reconstruction from an input image. In the MV-based category, Zero123 [liu2023zero] generates novel views from arbitrary perspectives, which can be further combined with SDS loss [poole2022dreamfusion] to perform 3D reconstruction. Additionally, SyncDreamer [liu2023syncdreamer], Era3D [li2024era3d], and Wonder3D [long2024wonder3d] produce multi-view consistent images, using the Neus [wang2021neus] representation to reconstruct 3D geometry. In contrast, Unique3D [wu2024unique3d] directly optimizes mesh vertices for a reconstruction approach. For feed-forward methods, Point-E [nichol2022point] and Shap-E [jun2023shap] are 3D generative models trained on a large internal OpenAI 3D dataset, both of which are able to convert a single-view image into a point cloud or an implicit representation. One-2-3-45 [liu2023one], CRM [wang2024crm] and InstantMesh [xu2024instantmesh] first generate multi-view results and then leverage a sparse-view feed-forward reconstruction model to deliver a textured mesh using the representation of triplane or implicit SDF representation.
@@ V-C Evaluation Protocol
Evaluation Datasets. Following prior research [liu2023zero, liu2023syncdreamer], we adopt the Google Scanned Object dataset [downs2022google] for our evaluation, which includes a wide variety of common everyday objects. Our evaluation dataset aligns with that of SyncDreamer [liu2023syncdreamer], comprising 30 objects spanning from everyday items to animals. For each object in the evaluation set, we render an image with a size of 256×256, which serves as the input. Additionally, we include in-the-wild images with diverse styles, collected from the internet or generated using text-to-image models, in our evaluation.
Metrics. To evaluate the quality of the single-view reconstructions, we adopt two metrics: Chamfer Distances (CD) and Volume IoU between ground-truth shapes and reconstructed shapes on both orthogonal and perspective settings of our method. Since different methods adopt various canonical systems, we first align the generated shapes to the ground-truth shapes before calculating the two metrics. Moreover, we adopt the metrics PSNR, SSIM [wang2004image] and LPIPS [zhang2018unreasonable] for evaluating the generated color images.
@@ V-D Single View Reconstruction
We compare the quality of the reconstructions with SOTA methods, the quantitative results are summarized in Table I, and the qualitative comparisons are presented in Fig. 6. Rather than modeling multi-view images and normal maps distribution jointly, Unique3D [wu2024unique3d] sequentially generate multi-view colors and normal maps. This exacerbates the geometric inconsistency since the stage 2 model has a huge domain gap in the inputs during training and in the inference period(see Section V-F for a more detailed discussion). As a result, vertices are updated by unstable and incorrect gradients during the optimization process, destabilizing the process and ultimately leading to a reconstructed object with incorrect shapes. Leveraging the generalization performance of the multi-view diffusion model and sparse-view reconstruction model, InstantMesh [xu2024instantmesh] produces results with reasonable overall geometry. However, due to the limited resolution of the triplane representation, the generated geometry and textures lack sufficient detail and exhibit artifacts aligned perpendicular to the coordinate axes. Wonder3D [long2024wonder3d] takes relatively low-resolution multi-view images and normal maps to reconstruct 3D objects using implicit SDF representation without an enhancement process, it also struggles to obtain fine geometric and texture details.
Thanks to our multi-stage training scheme and cascaded 3D mesh extraction, our method produces textured meshes with a well-defined overall shape and superior geometric and textural quality compared to baseline approaches. Furthermore, geometric quantization comparisons on the GSO dataset [downs2022google] highlight the generalization ability and effectiveness of our method (see Table  I).
@@ V-E Novel View Synthesis
We evaluate the quality of novel view synthesis for different multi-view-based methods. The quantitative results are presented in Table II, and the qualitative results can be found in Figure 5. SyncDreamer [liu2023zero] introduces a volume attention scheme to enhance the consistency of multi-view images, their model is sensitive to the elevation degrees of the input images and tends to produce unreasonable results. Unique3D [wu2024unique3d] stitching together the multiple views into a single image to generate multi-view results. This method severely limits the number of viewpoints it can produce (up to four). Moreover, generating images and normal maps in a sequential manner limits the generalizability and cross-view consistency of the model. Era3D [li2024era3d] proposes row-wise attention to ensure multi-view consistency and estimates the input camera focal length to normalize the output results into an orthogonal coordinate system. Although it generates results with a higher resolution, it is difficult to stably output reasonable multi-view results for in-the-wild images.
@@ V-F Discussions
Figure 11: The qualitative comparisons of reconstruction method between the geometry fusion algorithm proposed by Wonder3D [long2024wonder3d] and cascaded 3D mesh extraction method. Our approach has significant advantages in both geometric and textural details.
@@ Acknowledgment
This work was supported by Project NSFC 62501261. Additionally, this research was also supported by the Lhasa Science and TechnologyProgram Project titled ”Empowering Grassroots Social Governance Modernization through 5G and Multimodal Big Data Technologies”(Project No.LSKJ-202402).
VII Biography Section
[Uncaptioned image] Yuxiao Yang received his B.S. degree in Information Engineering from Zhejiang University, Hangzhou, China, in 2023. He is currently pursuing a master’s degree at the Shenzhen International Graduate School, Tsinghua University. His research interests include 3D generation and video generation.
[Uncaptioned image] Xiaoxiao Long is an Tenure-Track Associate Professor at School of Intelligence Science and Technology, Nanjing University. He obtained his Ph.D degree at the University of Hong Kong in 2023. His research includes 3D vision, computer graphics and robotic AI.
[Uncaptioned image] Zhiyang Dou is a Ph.D. candidate in the Computer Graphics Group at the University of Hong Kong, supervised by Prof. Wenping Wang and Prof. Taku Komura. He received his B. Eng. degree with honors from Shandong University, advised by Prof. Shiqing Xin. He is currently a visiting Ph.D. student in the Department of Computer and Information Science at the University of Pennsylvania, working at the Graphics Lab and GRASP Lab.
[Uncaptioned image] Cheng Lin received his Ph.D. from The University of Hong Kong (HKU), advised by Prof. Wenping Wang. Now he is an assistant professor at Macau University of Science and Technology. He visited the Visual Computing Group at Technical University of Munich (TUM), advised by Prof. Matthias Nießner. Before that, he completed his B.E. degree at Shandong University. His research interests include geometric modeling, 3D vision, shape analysis, and computer graphics.
[Uncaptioned image] Yuan Liu is visiting NTU as a PostDoc working with Prof. Ziwei Liu. He obtained his PhD degree at HKU advised by Prof. Wenping Wang and Prof. Taku Komura. His research mainly concentrates on 3D vision and graphics. He currently work on topics about 3D AIGC including neural rendering, neural representations, and 3D generative models.
[Uncaptioned image] Qingsong Yan is currently a Ph.D. student at the School of Geodesy and Geomatics, Wuhan University. He received the B.Eng. degree and MSc degree from Wuhan University in 2016 and 2019, respectively. He is majoring in omnidirectional computer vision, 3D reconstruction, multiview-stereo, and neural implicit representations.
[Uncaptioned image] Yuexin Ma received the PhD degree from the Department of Computer Science at University of Hong Kong in 2019. Now she is an assistant professor in ShanghaiTech University, China, leading 4DV Lab. Her research interests lie on the computer vision and deep learning. Particularly she is interested in 3D scene understanding, multi-modal perception, human-machine cooperation, and autonomous driving. She has published dozens of top-tier journal and conference papers in computer vision and robotics, including Science Robotics, TPAMI, IJCV, CVPR, ECCV, AAAI, TOG, TVCG, SIGGRAPH, ICRA, etc.
[Uncaptioned image] Haoqian Wang (Member, IEEE) received the PhD degree from the Harbin Institute of Technology, Harbin, in 2005. He was a post-doctoral fellow with Tsinghua University, Beijing, China, from 2005 to 2007. He has been a faculty member with the Shenzhen International Graduate School, Tsinghua University, Shenzhen, China, since 2008, where he has also been an associate professor and a professor since 2011 and the director of the Shenzhen Institute of Future Media Technology. His current research interests include Stereo Vision and deep learning.
[Uncaptioned image] Zhiqiang Wu is the Brage Golding Distinguished Professor of Electrical Engineering of Wright State University. He received his B.S. from Beijing University of Posts and Telecommunications in 1993, his M.S. from Peking University in 1996, and his Ph.D. from Colorado State University in 2002. He teaches courses in communication, signal processing and electronic warfare. His research interests include artificial intelligence, communication, signal processing, electronic warfare, and big data. His research has been funded by NSF, AFRL, AFOSR, ONR, NASA, DOE, and OFRN. He has published more than 200 papers in journals and conferences.
[Uncaptioned image] Wei Yin is a staff research scientist at Horizon Robotics. Before that, he received Ph.D. degree at Australian Institure of Machine Learning (AIML), University of Adelaide. His research interests include the world model, 3D reconstruction, and depth estimation.
@@ Appendix A Coordinate System
In practice, the target object is assumed to be placed along the gravity direction. 1) Canonical coordinate system. Some prior works (e.g. MVDream [shi2023mvdream] and SyncDreamer [liu2023syncdreamer]) adopt a shared canonical system for all objects, whose axis $Z_{c}$ shares the same direction with gravity (Fig. 15 (a)). 2) Input view related system. Wonder3D and Wonder3D++ adopt an independent coordinate system for each object that is related to the input view. Its $Z_{v}$ and $X_{v}$ axes are aligned with the UV dimension of 2D input image space, and its $Y_{v}$ axis is vertical to the 2D image plane and passes through the center of ROI (Region of Interests) (Fig. 15 (b)). 3) Camera poses. Wonder3D and Wonder3D++ outputs 6 views $\{v_{i},i=0,...,5\}$ that are sampled at the $X_{v}OY_{v}$ plane of the input-view related system with a fixed radius, where the front view $v_{0}$ is initialized as input view and the other views are sampled with pre-defined azimuth degrees (see Fig. 15 (b)).